OpenStudy (anonymous):

What identity is used in this function, and how do I solve it: sin^2[1/2cos^-1(4/5)] My book has no examples of this and I am lost

OpenStudy (freckles):

let u=arccos(4/5) then cos(u)=4/5  find the opp also but use the triangle here to find sin^2(1/2*u)

OpenStudy (anonymous):

the opp is: 3

OpenStudy (freckles):

recall \[\sin^2(\frac{u}{2})=\frac{1}{2}(1-\cos(u))\]

OpenStudy (freckles):

and right so that makes cos(u)=3/5 just plug in

OpenStudy (freckles):

and evaluate

OpenStudy (freckles):

no I'm dumb

OpenStudy (freckles):

cos(u)=4/5

OpenStudy (freckles):

you didn't have to find the opp side don't know what I was thinking there

OpenStudy (anonymous):

so it involves the half angle formula. Therefore i would take the angle of cos(u)=4/5 and evaluate it at \[(1/2)(1-\cos(4/5)\]

OpenStudy (anonymous):

forgot an extra ) on my equation

OpenStudy (freckles):

well you mean cos(u) is 4/5 not cos(u) is cos(4/5)

OpenStudy (freckles):

\[\frac{1}{2}(1-\cos(u)) \\ \text{ and we have} \cos(u)=\frac{4}{5} \\ \text{ so we have } \\ \frac{1}{2}(1-\frac{4}{5})\]

OpenStudy (anonymous):

oh yes. so: \[(1/2)(1-4/5)\]

OpenStudy (freckles):

little arithmetic and you are done

OpenStudy (anonymous):

That was much more simple than I believed it would be. Thank you so much!

OpenStudy (freckles):

np